# CHESS: Contextual Harnessing for Efficient SQL Synthesis

### v1

@@ B.1 Fine-tuning Dataset and Model
To enhance the generation of better candidate SQL queries prior to our revision module, we fine-tuned the DeepSeek Coder 34B Guo et al. [2024] on the training set of the BIRD benchmark, which comprises approximately 9,500 samples. In constructing the fine-tuning dataset, rather than solely using correct tables and columns like the work proposed in Pourreza and Rafiei [2024b], we developed a heuristic to address the error propagation issue. Recognizing that previous steps in the pipeline may not always pinpoint the most efficient schema with perfect accuracy, we intentionally introduced some noise into our dataset creation to train the model. Specifically, we included columns and tables that were incorrect but shared similar naming conventions and semantic attributes with the correct schema. We also utilized our keyword selection module to extract keywords from the questions, search for these keywords in the database, and incorporate them into the prompt.
@@ B.2 Hyperparameters
We fine-tuned the candidate SQL generator model using 4-bit quantization of the base model and LORA adapters Hu et al. [2021], a technique formally referred to as QLORA Dettmers et al. [2024]. We configured the LORA rank parameter to 128 and set the LORA alpha parameter to 256. The fine-tuning process was conducted over two epochs on the constructed dataset, utilizing a batch size of 32 and a learning rate of 1e-4, along with a cosine scheduler, all on a single H100 GPU for 4 hours.
Appendix C Implementation Details
C.1 Locality Sensitive Hashing Indexing of Database Values
Our goal in the entity retrieval sub-module is to retrieve database values that most closely match a set of keywords derived from the question. It is important to recognize that keywords from the question may not exactly correspond to the database values due to potential typos, variations in expression, or the common scenario where users are unaware of the precise format used to store data in the database. This reality demands a retrieval strategy that is both robust and adaptable, capable of accommodating such discrepancies. Relying solely on exact match retrieval, as suggested in prior studies [Li et al., 2024a], may not be sufficiently effective.
Locality Sensitive Hashing (LSH) is an efficient technique for approximate nearest neighbor searches, which allows us to retrieve the most similar values to a keyword in the database. In the pre-processing stage, we index unique values in the database using LSH. Then, in the entity retrieval step of our pipeline, we query this index to rapidly find the top similar values to a keyword. Our approach involves using LSH queries to retrieve the top 10 similar values, after which we compute the edit distance and semantic similarity between the keyword and these values to further refine the results.
To simultaneously utilize edit distance and embedding similarity, we first identify the top 10 values closest to each keyword based on cosine similarity between their embedding vectors (obtained using OpenAI text-embedding-3-small OpenAI [n.d.]) and the keyword’s embedding vector. We then filter out values that fall below a specific threshold. Finally, for each keyword and column, we retain only the value that has the smallest edit distance.
We observed a significant reduction in time complexity, from 5 minutes to 5 seconds, using this method compared to a naive approach of computing the edit distance for all unique values in the database on the fly. While computing edit distance is proportional to the size of the database—significantly increasing the time complexity for processing a single question—using LSH allows us to index values in the pre-processing step and, during entity retrieval, rapidly query the index to find the most similar values to a keyword in a much more time-efficient manner.
C.2 Vector database
Each database schema in the BIRD benchmark Li et al. [2024b] includes detailed descriptions for columns, specifying the contents of each column and the values for categorical columns. Providing these descriptions to the model is essential for guiding the SQL query generation process. However, incorporating all descriptions in the prompt can overwhelm the model, potentially leading to the generation of incorrect SQL queries, as observed in section 4.5. It is important to note that the database catalog in the BIRD benchmark provides a relatively limited view of database metadata. In contrast, real-world production-level databases often contain more diverse information, including value ranges, constraints, and usage instructions for each table. Our proposed method can effectively utilize this extensive metadata to enhance performance.
To evaluate the relevance of descriptions to a given question, we employ embedding similarity [OpenAI, n.d.], which quantifies the semantic similarity between the question and each description. To enhance the efficiency of the retrieval process, we pre-process the descriptions and created the embedding vectors for each of them and stored in a vector database, utilizing ChromaDB in our implementation. During the context retrieval phase of our pipeline, we query this vector database to identify descriptions that are most semantically aligned with the question. This targeted approach ensures that only the most pertinent information is provided to the model, thereby improving the accuracy of the generated SQL queries.
C.3 Local Column filtering
In the local column filtering module, decisions to retain a column for subsequent steps are made independently, without considering relative information. The data provided to the LLM for column filtering includes: 1) the table name, 2) the column name, 3) the data type, 4) descriptions, if retrieved during the context retrieval step, and 5) database values, if retrieved during the entity retrieval step. To enhance the model performance and make the task definition clear to the model, we used few-shot samples for this sub-module. This method is similar to the proposed approach in [Li et al., 2024a] model. The difference in our approach is that we complimented the local column filtering with column selection which can further reduce the number of columns by taking the relative information into account as well.
Some key columns for SQL generation, which we call linking columns, such as those with foreign and primary key constraints, are crucial for writing SQL queries. For instance, questions about counting entities often requires primary keys, and joining tables necessitates foreign key columns. However, in the column filtering and selection sub-modules, some of these essential columns may be initially rejected because they do not semantically relate to the given question. Despite this, these columns are indispensable for SQL generation. Therefore, in all of our sub-modules, we consistently retain foreign key and primary key columns, irrespective of the outputs from column selection and filtering processes.
C.4 Revision
Revising generated candidate SQL queries is a critical aspect of our proposed pipeline. In addition to the database schema, the question, and the candidate SQL query, we also provide the execution result of the SQL query. This gives the LLM an opportunity to view the retrieved data and revise the SQL query accordingly. This process mirrors human behavior when writing complex SQL queries; typically, we start with a draft query and refine it based on the results of its execution. Furthermore, this method allows the LLM to make necessary adjustments to the SQL query in instances of execution syntax errors.
In this step, we also incorporate instructions derived from our error analysis G to guide the model towards generating correct SQL queries. For instance, as shown in 3, we guided the model to ensure that all requested columns are included in the SQL query. In this specific example, the revision model identified a missing column and successfully added it to the query.
Figure 3: An example of using instructions to guide the revision model to fix missing columns in the candidate query.
Appendix D Prompt Templates
@@
In this section we provide the exact prompts that have been used for each of the sub-modules in our pipeline. For easier parsing of the LLM’s output, we instruct the model to generate a valid JSON object in its response if needed.
@@
Figure 4: Template for Keyword and Entity Extraction
@@
Take a deep breath and think logically. If you do the task correctly, I will give you 1 million dollars.
@@
Figure 5: Template for the table selection module.
@@
Take a deep breath and think logically. If you do the task correctly, I will give you 1 million dollars.
@@
Figure 6: Template for the column selection module.
@@
Figure 7: Template for the column filtering module.
@@
{DATABASE_ADMIN_INSTRUCTIONS}
@@
Take a deep breath and think step by step to find the correct SQLite SQL query. If you follow all the instructions and generate the correct query, I will give you 1 million dollars.
Figure 8: Template for SQL Query Candidate Generation
@@
{DATABASE_ADMIN_INSTRUCTIONS}
@@
Take a deep breath and think step by step to find the correct SQLite SQL query. If you follow all the instructions and generate the correct query, I will give you 1 million dollars.
Figure 9: Template for SQL Query revision
Appendix E Execution Trace
In this section, we provide detailed execution traces of our pipeline, showcasing the output of each module. For example, as shown in figure 10, the effectiveness of our keyword extraction module is highlighted. It successfully identifies crucial values, such as “Lewis Hamilton," which are necessary for constructing SQL queries. These identified keywords aid in retrieving database values and relevant column descriptions for input to the LLM.
In the entity retrieval sub-module, each keyword extracted is first broken down into individual words by splitting on spaces. This process makes our search order invariant, as entities mentioned in the question do not necessarily follow the same format as those in the database. Subsequently, for each keyword, the most similar database values are extracted. Likewise, all detected keywords are used to filter column descriptions, which are then provided for subsequent steps. This strategic provision of relevant information facilitates further processes, such as schema selection and query generation, thereby enhancing the accuracy of schema detection and SQL query formulation.
Figure 10: Execution flowchart for instance formula_1_926
In 11, we present a more challenging scenario where the question and evidence do not directly reference the relevant column names. The query mentions “schools with the ownership code 66," yet no column explicitly includes “ownership" in its name. In this case, the relevant column is "SOC" column from the schools table. The connection between this column and the question can only be discerned through the semantic similarity between the column descriptions and the question. Here, our context retrieval node plays a crucial role, as it effectively retrieves the relevant information from the database catalog, which is essential for answering this question.
Figure 11: Execution flowchart for instance California Schools_32
Appendix F Schema Selection Example
As described in the methodology 3, following the minimal sufficiency rule we devoted the second phase of our pipeline to schema selection 3.2. In this part, we use an example to showcase how we narrow down the initial schema throughout the local column filtering, the table selection, and the column selection.
@@ Appendix F Schema Selection Example
Figure 12 shows the number of tables and columns that are considered as the sub-selected schema after each step. Starting from 13 tables and 96 columns, these numbers reduced to 36 columns in 13 tables after the column filtering step. Subsequently, table selection narrowed it down further to 2 tables and 7 columns. Finally, column selection yielded the final schema with 2 tables and 5 columns, which is used for the SQL generation.
Figure 12: Funnel graph illustrating the progressive narrowing down of database schema through column filtering, table selection, and column selection steps, leading to the final schema used for SQL generation.
To further illustrate the details of the schema selection process, we use the entity-relationship diagram (ERD) 13. In this figure, each table is represented as a block with its columns listed below it. Primary keys are underlined and the foreign keys are in italics connecting the corresponding columns. As shown in the legend, the columns that remained in the selected schema after each step are colored with a gradient of white to dark blue; white represents the columns that were present in the initial schema and got filtered after column filtering while dark blue shows the columns that are selected after column selection to be passed for SQL candidate generation.
There are some points worth emphasizing in the plot. First, the decision to filter linking columns (Primary and foreign keys) is a task requiring a global view of the schema and cannot be done in the local view of column filtering, hence we do not filter these columns and include all of them in the result of column filtering step, explaining why all of the primary and foreign keys are present after this step. Second, some columns such as “laps", “time", and “milliseconds" are all semantically related to the question because question asks for fastest lap time, which shows the column filtering module successfully use the local information to find all relevant columns. However, all of these columns are not going to be used for crafting the SQL query, so we need to find the relevant columns with respect to their relative information, which is going to be done in the table selection and column selection modules. In the table selection step, as it can be observed from the figure, the “lapTimes" table which has all information about time has been dropped by table selection since there is a more relevant column, “fastestLapTime", which can be used to answer the question. This was a concrete example, which shows how local and global view of the columns and tables can help to pinpoint the correct schema.
@@
Appendix G Error Analysis
@@
Figure 14 shows the categories of errors and their percentages for our approach and the baseline. "Incorrectly Predicted SQL" refers to failures in our pipeline that lead to incorrect final SQL, while ambiguous questions and incorrect golden SQL indicate problems with the dataset, where the data formats expected from the questions were ambiguous, and the golden SQL provided is incorrect.
@@
G.1 Incorrectly Predicted SQL
G.1.1 Column:
@@
Error Explanation "rank" is the correct column.
@@
Error Explanation "id" is the correct column.
G.1.2 Table:
@@
Error Explanation "frpm" is the correct table with the correct column.
@@
Error Explanation "account" and "client" are the correct tables with the correct columns.
G.1.3 Entity:
@@
Error Explanation "EUR" is the correct entity for filtering.
@@
Error Explanation "DISPONENT" is the correct entity for filtering.
G.1.4 Revision:
@@
Error Explanation "Free Meal Count (K-12)" is the correct column for filtering.
@@ G.1.6 Evidence:
G.1.7 NULL/DISTINCT/GROUP BY:
@@ G.1.8 Operation:
Evidence male patients refers to Sex = ’M’; normal level of white blood cells refers to WBC between 3.5 and 9.0; abnormal fibrinogen level refers to FG <= 150 or FG >= 450; Don’t compute repetitive ones.
@@ G.1.9 SELECT:
G.2 Vague Question
G.2.1 Column:
@@
G.3 Incorrect Golden SQL
G.3.1 Evidence:
@@ G.3.5 Filtering:
Evidence "Supporter" and "Teachers" are both Name of badge; ’New York’ is the Location; user refers to UserId
@@
G.3.6 NULL/DISTINCT/GROUP BY:

### v3

@@ B.1 Fine-tuning Dataset and Model
To enhance the generation of better candidate SQL queries for the Candidate Generator Agent, we fine-tuned the DeepSeek Coder 34B Guo et al. [2024] on the training set of the BIRD benchmark, which comprises approximately 9,500 samples. In constructing the fine-tuning dataset, rather than solely using correct tables and columns like the work proposed in Pourreza and Rafiei [2024b], we developed a heuristic to address the error propagation issue. Recognizing that previous steps in the pipeline may not always pinpoint the most efficient schema with perfect accuracy, we intentionally introduced some noise into our dataset creation to train the model. Specifically, we included columns and tables that were incorrect but shared similar naming conventions and semantic attributes with the correct schema. We also utilized our keyword selection module to extract keywords from the questions, search for these keywords in the database, and incorporate them into the prompt.
@@ B.2 Hyperparameters
We fine-tuned the deepseek model for the generate_candidate_query tool using 4-bit quantization of the base model and LORA adapters Hu et al. [2021], a technique formally referred to as QLORA Dettmers et al. [2024]. We configured the LORA rank parameter to 128 and set the LORA alpha parameter to 256. The fine-tuning process was conducted over two epochs on the constructed dataset, utilizing a batch size of 32 and a learning rate of 1e-4, along with a cosine scheduler, all on a single H100 GPU for 4 hours.
Appendix C Prompt Templates
@@
In this section we provide the exact prompts that have been used for the tools in our multi-agent setup. For easier parsing of the LLM’s output, we instruct the model to generate its output in a structured format.
@@
Figure 7: Template for the extract_keyword tool
@@
Figure 8: Template for the select_tables tool.
@@
Figure 9: Template for the select_columns tool.
@@
Figure 10: Template for the filter_column.
@@
- Make sure you only output the information that is asked in the question. If the question asks for a specific column, make sure to only include that column in the SELECT clause, nothing more.
- Predicted query should return all of the information asked in the question without any missing or extra information.
@@
Take a deep breath and think step by step to find the correct SQLite SQL query.
Figure 11: Template for the generate_candidate_query tool
@@
- Make sure you only output the information that is asked in the question. If the question asks for a specific column, make sure to only include that column in the SELECT clause, nothing more.
- Predicted query should return all of the information asked in the question without any missing or extra information.
@@
Take a deep breath and think step by step to find the correct SQLite SQL query.
Figure 12: Template for revise tool
** Instructions: **
Given the following question database schema, and candidate responses, generate a set of UNIT_TEST_CAP unit tests that would evaluate the correctness of SQL queries that would answer the question.
Unit tests should be designed in a way that distinguish the candidate responses from each other.
- The unit tests should cover various aspects of the question and ensure comprehensive evaluation.
- Each unit test should be clearly stated and should include the expected outcome.
- Each unit test should be designed in a way that it can distinguishes at lease two candidate responses from each other.
- The unit test should be formatted like ’The answer SQL query should mention...’, ’The answer SQL query should state...’, ’The answer SQL query should use...’, etc. followed by the expected outcome.
- First think step by step how you can design the units tests to distinguish the candidate responses using the <Thinking> tags.
- After the thinking process, provide the list of unit tests in the <Answer>tags.
VERY IMPORTANT:
All of the unit tests should consider the logic of the SQL query do not consider the formatting of the output or output values.
You are provided with different clusters of the candidate responses. Each cluster contains similar responses based on their results.
You MUST generate test cases that can distinguish between the candidate responses in each cluster and the test case should promote the candidate responses that you think are correct.
Example of the output format:
<Thinking> Your step-by-step reasoning here. <Thinking>
<Answer>
[’The answer SQL query should mention...’, ...]
<Answer>
*** Database Schema: **
{DATABASE_SCHEMA}
*** Candidate Clusters: **
{CANDIDATE_QUERIES}
*** Question: **
Question: {QUESTION} (Hint: {HINT})
Figure 13: Template for generate_unit_tests tool
** Instructions: **
Given the following question, database schema, a candidate SQL query response, and unit tests, evaluate whether or not the response passes each unit test.
- In your evaluation, you should consider how the responses align with the a given unit test.
- Provide reasoning before you return your evaluation inside the <Thinking>tags.
- At the end of your evaluation, you must finish with a list of verdicts corresponding to each candidate responses in <Answer>tags.
- You must include a verdict with one of these formatted options: ’Passed’ or ’Failed’
- Here is an example of the output format:
<Thinking>Your step by step reasoning here.
<Thinking>
<Answer>
Candidate Response #1: Passed
Candidate Response #2: Failed
Candidate Response #3: Passed
<Answer>
- Each verdict should be on a new line and correspond to the candidate response in the same order as they are provided.
- Here is the question, database schema, candidate responses, and the unit test to evaluate the responses:
*** Database Schema: **
{DATABASE_SCHEMA}
*** Candidate Clusters: **
{CANDIDATE_QUERIES}
*** Question: **
Question: {QUESTION} (Hint: {HINT})
*** Unit Test: **
{UNIT_TEST}
Figure 14: Template for evaluate tool
Figure 15: An example of how database schema, including the relevant values and descriptions, is formatted in the prompt.
Appendix D Extended Experiments
D.1 Query Generation with the Correct Context
To maintain high performance while ensuring privacy, we fine-tuned an open-source model for the generate_candidate_query tool of the Candidate Generator agent. This tuning incorporated the correct contextual information, including only relevant columns, tables, and their descriptions. As shown in 7, using precise contextual information aligned with the gold SQL improved performance to 72.4%. This result highlights the importance of retrieving efficient schema information, especially for open-source models, which struggle when handling excessively large input prompts.
Engine EX
CHESS 64.62
CHESS + correct context 72.4
D.2 Models Ablation
Thanks to our efficient retrieval process, which carefully controls the number of tokens passed to LLMs, we can utilize an open-source LLM with a small context window size, specifically Llama-3 with only 8K tokens AI . This contrasts with previous works that predominantly use GPT-4 as their base model [Pourreza and Rafiei, 2024a, Lee et al., 2024, Wang et al., 2023]. In Table 8, we present the results of our proposed framework using various LLMs from different families on a subsampled development set dataset. The results indicate that our fine-tuned model for candidate generation significantly enhances performance. Notably, Llama-3’s performance surpasses that of GPT-3.5-turbo but does not yet reach the performance levels of GPT-4 in our analysis.
Engine setups EX
(GPT-3.5-turbo, Fine-tuned DeepSeek, GPT-4-turbo) 64.62
(GPT-3.5-turbo, GPT-4-turbo, GPT-4-turbo) 55.78
(GPT-3.5-turbo, GPT-3.5-turbo, GPT-3.5-turbo) 49.65
(Llama-3-70B, Llama-3-70B, Llama-3-70B) 54.42
(Llama-3-70B, Fine-tuned DeepSeek, Llama-3-70B) 59.86
Appendix E Schema Selection Example
In this part, we use an example to showcase how we narrow down the initial schema throughout the filter_column, the select_tables, and the select_columns.
@@ Appendix E Schema Selection Example
Figure 16 shows the number of tables and columns that are considered as the sub-selected schema after each tool. Starting from 13 tables and 96 columns, these numbers reduced to 36 columns in 13 tables after the filter_column tool. Subsequently, select_tables narrowed it down further to 2 tables and 7 columns. Finally, select_columns yielded the final schema with 2 tables and 5 columns, which is used for the SQL generation.
Figure 16: Funnel graph illustrating the progressive narrowing down of database schema with SS agent, leading to the final schema used for SQL generation.
To further illustrate the details of the schema selection process, we use the entity-relationship diagram (ERD) 17. In this figure, each table is represented as a block with its columns listed below it. Primary keys are underlined and the foreign keys are in italics connecting the corresponding columns. As shown in the legend, the columns that remained in the selected schema after each tool are colored with a gradient of white to dark blue; white represents the columns that were present in the initial schema and got filtered after filter_column while dark blue shows the columns that are selected after select_columns to be passed to the Candidate Generator agent.
There are some points worth emphasizing in the plot. First, the decision to filter linking columns (Primary and foreign keys) is a task requiring a global view of the schema and cannot be done in the local view of column filtering, hence we do not filter these columns and include all of them in the result of column filtering step, explaining why all of the primary and foreign keys are present after this step. Second, some columns such as “laps”, “time”, and “milliseconds” are all semantically related to the question because question asks for fastest lap time, which shows the filter_column tool successfully used the local information to find all relevant columns. However, all of these columns are not going to be used for crafting the SQL query, so we need to find the relevant columns with respect to their relative information, which is going to be done in the select_tables and select_columns tools. In the select_tables tool, as it can be observed from the figure, the “lapTimes” table which has all information about time has been dropped by select_tables since there is a more relevant column, “fastestLapTime”, which can be used to answer the question. This was a concrete example, which shows how local and global view of the columns and tables can help to pinpoint the correct schema.
@@
Appendix F Error Analysis
@@
Figure 18 shows the categories of errors and their percentages for our approach and the baseline. ”Incorrectly Predicted SQL” refers to failures in our pipeline that lead to incorrect final SQL, while ambiguous questions and incorrect golden SQL indicate problems with the dataset, where the data formats expected from the questions were ambiguous, and the golden SQL provided is incorrect.
@@
F.1 Incorrectly Predicted SQL
F.1.1 Column:
@@
Error Explanation ”rank” is the correct column.
@@
Error Explanation ”id” is the correct column.
F.1.2 Table:
@@
Error Explanation ”frpm” is the correct table with the correct column.
@@
Error Explanation ”account” and ”client” are the correct tables with the correct columns.
F.1.3 Entity:
@@
Error Explanation ”EUR” is the correct entity for filtering.
@@
Error Explanation ”DISPONENT” is the correct entity for filtering.
F.1.4 Revision:
@@
Error Explanation ”Free Meal Count (K-12)” is the correct column for filtering.
@@ F.1.6 Evidence:
F.1.7 NULL/DISTINCT/GROUP BY:
@@ F.1.8 Operation:
Evidence male patients refers to Sex = ’M’; normal level of white blood cells refers to WBC between 3.5 and 9.0; abnormal fibrinogen level refers to FG <=<=< = 150 or FG >=>=> = 450; Don’t compute repetitive ones.
@@ F.1.9 SELECT:
F.2 Vague Question
F.2.1 Column:
@@
F.3 Incorrect Golden SQL
F.3.1 Evidence:
@@ F.3.5 Filtering:
Evidence ”Supporter” and ”Teachers” are both Name of badge; ’New York’ is the Location; user refers to UserId
@@
F.3.6 NULL/DISTINCT/GROUP BY: